OpenStudy (anonymous):

HELP, PLEAASSEEE???? If n = 4, all of the following expressions are equivalent except _____. 16 1/n^2 1/16 n^-2

OpenStudy (shane_b):

\[\frac{1}{n^2}=\frac{1}{4^2}=\frac{1}{16}\]and \[n^{-2}=\frac{1}{n^2}\] So what answer do you think?

OpenStudy (anonymous):

1 over n^2?

OpenStudy (shane_b):

No...the last 3 options are equivalent...the first is not.

OpenStudy (anonymous):

so 16?

OpenStudy (shane_b):

Yes...since 1/16 (which the other options equate to) does not equal 16